The sum of three numbers is 8
.
The first number minus the second plus the third is 4
.
The first minus the third is 2

more than the second. Find the numbers.

Answer by Alan3354(69443)

The sum of three numbers is 8
.
The first number minus the second plus the third is 4
.
The first minus the third is 2

more than the second. Find the numbers.
----------
Call them x, y & z
x + y + z = 8
x - y + z = 4
x - z = y + 2 --> x - y - z = 2
--------
x + y + z = 8
x - y + z = 4
x - y - z = 2
-----------
Add eqns 1 & 3
x + y + z = 8
x - y - z = 2
-------------
2x = 10
x = 5
=====================
Add eqns 1 & 2
x + y + z = 8
x - y + z = 4
2x + 2z = 12
x + z = 6
z = 1
============
y = 2